# Approximate predictive control barrier function for discrete-time systems

### v1

@@ 5.1 Approximate PCBF
In the proof of Theorem 5.19, the error $|e_{h}(x)|$ can also be bounded by a class $\mathcal{K}$ function $\gamma$ and a constant term $\tilde{\epsilon}_{h}\leq\epsilon_{h}$, such that $|e_{h}(x)|\leq\gamma(x)+\tilde{\epsilon}_{h}$. If it holds that $\forall x\in\mathcal{D}_{\mathrm{PB}}$, $c_{\Delta}\Delta h_{\mathrm{PB}}(x)>2\gamma(x)$, the decrease condition is still satisfied with respect to the smaller constant $2\tilde{\epsilon}_{h}+2\epsilon_{h}$. This implies that approximation errors close to $\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$ have a bigger impact on the closed-loop stability than errors when $h_{\mathrm{PB}}$ is large, which can be leveraged when performing the approximation as discussed in Section 6. Furthermore, it can be shown that the approximation is additionally input to state stable with respect to an additional disturbance $w$ acting on system (1), following the results in Theorem 4.14.
5.2 Continuous extension of the PCBF
As the PCBF $h_{\mathrm{PB}}(x)=0$ for all $x\in\mathcal{S}_{\mathrm{PB}}$, errors in its approximation can lead to undesired conservativeness of the safety filter within $\mathcal{S}_{\mathrm{PB}}$ if $\hat{h}_{\mathrm{PB}}(x)>0$, as potentially safe control inputs can be deemed unsafe. In order to avoid such unnecessary safety filter interventions, we propose a continuous extension of $h_{\mathrm{PB}}$ in $\mathcal{S}_{\mathrm{PB}}$, which is negative and therefore mitigates potential learning errors leading to the state being considered unsafe. Consider the function
$h(x):=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}\\ \bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}},\end{cases}$ (21)
where the extension on $\mathcal{S}_{\mathrm{PB}}$ is given by
$\bar{h}_{\mathrm{PB}}(x(k)):=\min_{u_{i|k},\xi_{i|k}}\;\;$ $\max_{i,j}\{[\xi_{i|k}]_{j},\xi_{N|k}\}$ (22a)
s.t. $\forall\,i=0,\dots,N-1,$
$x_{0|k}=x(k),$ (22b)
$x_{i+1|k}=f(x_{i|k},u_{i|k}),$ (22c)
$u_{i|k}\in\mathcal{U},$ (22d)
$x_{i|k}\in\overline{\mathcal{X}}_{i}(\xi_{i|k}),$ (22e)
$h_{f}(x_{N|k})\leq\xi_{N|k}.$ (22f)
Note that the slacks in (22e) and (22f) are not required to be positive, implying negativity of $\bar{h}_{\textup{PB}}$ when the state and terminal constraints are satisfied. We show that the function $h$ is continuous on $\mathcal{D}_{\mathrm{PB}}$ and a control barrier function according to Definition 1.
Theorem 5.22.
Consider the function $h:\mathcal{D}_{\mathrm{PB}}\rightarrow\mathbb{R}$, with
$h(x)=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}},\\ \bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}}.\end{cases}$
It holds that $h$ is a CBF according to Definition 1 and the function $\tilde{h}(x)=\max\{0,h(x)\}=h_{\mathrm{PB}}(x)$ is a CBF according to Definition 4.
Proof 5.23.
We prove continuity of $h$ by proving continuity of $h_{\mathrm{PB}}$ on $\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, continuity of $\bar{h}_{\mathrm{PB}}$ on $\mathcal{S}_{\mathrm{PB}}$ and that $h_{\mathrm{PB}}(x)=\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$.
1) Continuity of $h_{\mathrm{PB}}$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$: In [12, Theorem III.6 b)], it is shown that $h_{\mathrm{PB}}$ is continuous for all $x\in\mathcal{D}_{\mathrm{PB}}$, which implies continuity of $h$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$.
2) Continuity of $\bar{h}_{\mathrm{PB}}$ for all $x\in\mathcal{S}_{\mathrm{PB}}$: We can show continuity of $\bar{h}_{\mathrm{PB}}$ by using similar arguments as in [12, Theorem III.6 b)]. For any $\epsilon>0$ such that $|\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})|<\epsilon$, with $x,\bar{x}\in\mathcal{S}_{\mathrm{PB}}$, we show that there exists a $\delta>0$, such that $\|x-\bar{x}\|<\delta$. Consider the optimal solution of (22) at $x$, i.e. $u_{i|k}^{*}(x)$ as a candidate solution in $\bar{x}$, i.e. $\bar{u}_{i|k}(\bar{x})=u^{*}_{i|k}$, which results in the state predictions $\bar{x}_{i|k}(\bar{x})$ with $\bar{x}_{0|k}=\bar{x}$. The resulting slacks are then given by $\bar{\xi}_{i|k}(\bar{x})=\min\{0,c(\bar{x}_{i|k})+\Delta_{i}\mathbf{1}\}$ and $\bar{\xi}_{N|k}=\min\{0,h_{f}(\bar{x}_{N|k}(\bar{x}))\}$. As the initial state and the corresponding inputs lie in a compact set, i.e. $\bar{x}\in\mathcal{S}_{\mathrm{PB}}$ and $\bar{u}_{i|k}(\bar{x})\in\mathcal{U}$, and the functions $f,c_{j}$ and $h_{f}$ are continuous, it holds that the candidate solution also lies within a compact set from [12, Lemma C.3] and from the Heine-Cantor theorem, it follows that $f,c_{j}$ and $h_{f}$ are uniformly continuous. As compositions of functions, the minimum and additions of uniform continuous functions all preserve uniform continuity, the state predictions, the candidate slack variables and the corresponding constructed candidate value function $\tilde{h}_{\mathrm{PB}}(\bar{x})=\max_{i,j}\{[\bar{\xi}_{i|k}(\bar{x})]_{j},\bar{\xi}_{N|k}(\bar{x})\}$ are uniformly continuous in the initial condition $\bar{x}$. From the fact that the optimal solution $\bar{h}_{\mathrm{PB}}(\bar{x})\leq\tilde{h}_{\mathrm{PB}}(\bar{x})$ due to the minimisation in (22), it holds that for any $\epsilon>0$, there exists a uniform $\delta>0$ with $\|x-\bar{x}\|<\delta$, such that $\bar{h}_{\mathrm{PB}}(\bar{x})-\bar{h}_{\mathrm{PB}}(x)\leq\tilde{h}_{\mathrm{PB}}(\bar{x})-\bar{h}_{\mathrm{PB}}(x)<\epsilon$. The same arguments hold when using the optimal solution at $\bar{x}$ in $x$, i.e. $\bar{u}_{i|k}(x)=u_{i|k}^{*}(\bar{x})$, such that there exists $\delta>0$ with $\|x-\bar{x}\|<\delta$, which implies that $\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})\leq\tilde{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})<\epsilon$, which in turn implies that for any $\epsilon>0$, there exists $\delta>0$ with $\|x-\bar{x}\|<\delta$, which implies that $|\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})|<\epsilon$.
3) Equality $\bar{h}_{\mathrm{PB}}(x)=h_{\mathrm{PB}}(x)$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$: By definition, we have $h_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$. In order to show that $\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$ we first consider an optimal input sequence $\tilde{u}_{i|k}^{*}(\tilde{x})$ of (22) for any $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$. Given this optimal input sequence, we showed in 2)2)2 ) that the corresponding slack variables $\tilde{\xi}_{i|k}(\tilde{x})$ and $\tilde{\xi}_{N|k}(\tilde{x})$ and the optimal value function $\bar{h}_{\mathrm{PB}}(\tilde{x})$ are uniformly continuous in $\tilde{x}$. Furthermore, it holds that the candidate slacks $\tilde{\xi}_{i|k}(\tilde{x})\leq 0$ and $\tilde{\xi}_{N|k}(\tilde{x})\leq 0$ as any optimal inputs $u^{*}_{i|k}(\tilde{x})$ of (10) are feasible in (22) and $h_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}\subseteq\mathcal{S}_{\mathrm{PB}}$. As the candidate slacks are all non-positive, it follows that $\bar{h}_{\mathrm{PB}}(x)\leq 0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$. We can then show that $\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}\subseteq\mathcal{S}_{\mathrm{PB}}$ by contradiction. Suppose that for any $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, it holds that there exists a solution with all slack variables $\xi_{i|k}(\tilde{x})<0$ and $\xi_{N|k}(\tilde{x})<0$ in (22). Then for the corresponding optimal input sequence $\tilde{u}_{i|k}(\tilde{x})$, it holds that $\xi_{i|k}(\tilde{x})$ and $\xi_{N|k}(\tilde{x})$ are continuous in $\tilde{x}$. The continuity of the slack variables in $\tilde{x}$ implies that there exists an open ball of radius $\delta>0$, i.e. $B_{\mathcal{D}}^{\delta}(\tilde{x})=\{x\in\mathcal{D}\mid\|x-\tilde{x}\|<\delta\}$, for which it still holds that $\xi_{i|k}(\tilde{x})<0$ and $\xi_{N|k}(\tilde{x})<0$. As the optimal input $\tilde{u}_{i|k}(\tilde{x})$ is a feasible input for (10), it implies that $h_{\mathrm{PB}}(x)=0$ for all $x\in B_{\mathcal{D}}^{\delta}(\tilde{x})$, which implies that $B_{\mathcal{D}}^{\delta}(\tilde{x})\subseteq\mathcal{S}_{\mathrm{PB}}$. This fact however contradicts with the initial assumption that $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, as by definition any open neighbourhood of $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$ must contain at least one point in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$. By contradiction, it therefore holds that any optimal solution of (22) has at least one slack $[\tilde{\xi}^{*}_{i|k}(\tilde{x})]_{j}$ or $\tilde{\xi}^{*}_{N|k}(\tilde{x})$ which is 0 and all slacks are non-positive, which implies that $\bar{h}_{\mathrm{PB}}(\tilde{x})=\max_{i,j}\{[\tilde{\xi}^{*}_{i|k}(\tilde{x})]_{j},\tilde{\xi}^{*}_{N|k}(\tilde{x})\}=0$.
As the function $h$ is a piece-wise function which is continuous on both of its subdomains $\mathcal{S}_{\mathrm{PB}}$ and $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and $\bar{h}_{\mathrm{PB}}(x)=h_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, it holds that $h$ is continuous on $\mathcal{D}_{\mathrm{PB}}$.
Remark 5.24.
Note that the function
$h(x):=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}},\\ c_{h}\bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}}.\end{cases}$ (23)
In principle, this continuous extension requires solving two optimisation problems for each sample $x\in\mathcal{S}_{\mathrm{PB}}$, i.e., (10) to determine if $x\in\mathcal{S}_{\mathrm{PB}}$ or $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and (22) in case $x\in\mathcal{S}_{\mathrm{PB}}$. However, a geometric sampler such as used in [41] can be used, which performs a line search over given seed and target points. By using such a geometric sampler, information of the previous solution can be used to solve either (10) or (22) based on which problem was previously solved. It follows that two optimisation problems only need to be solved if the samples on the line cross over from $\mathcal{S}_{\mathrm{PB}}$ into $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ or vice-versa. Additionally, this geometric sampler benefits from being able to warmstart the solver using the previous optimal solution $\mathbf{o}=\{x^{*}_{i},u^{*}_{i},\xi^{*}_{i},\bm{\nu}^{*},\bm{\lambda}^{*}\}$, where $\bm{\nu}^{*}$ and $\bm{\lambda}^{*}$ denote the optimal dual variables. Given that the optimal solution is shown to be continuous in $x$ in [12, Theorem III.6] and Theorem 5.22, the previously computed optimal solution will be close to the solution of the next iteration in the line search. This allows the computation time during sampling to be significantly reduced as shown in [41]. Finally, the random walk can be parallelised by performing line search for different goal points given a random starting seed. The resulting sampling algorithm is then given by Algorithm 2 in the Appendix.
@@ 6.1 Linear system
$x(k+1)$ $=Ax(k)+Bu(k)$ (24)
$=\begin{bmatrix}1&T_{s}\\ T_{s}\frac{g}{l}&1-T_{s}c\end{bmatrix}x(k)+\begin{bmatrix}0\\ 5T_{s}\end{bmatrix}u(k),$
where the states are given by the angle and angular velocity, i.e., $x=\left[\psi\;\;\omega\right]^{\top}\in\mathbb{R}^{2}$, the input $u\in\mathbb{R}$ and the dynamics are defined through the sampling time $T_{s}=0.1$, the damping coefficient is $c=1$, the length $l=1.3$ and $g=9.81$. The system is subject to polytopic state and input constraints $x(k)\in\mathcal{X}=\{x\in\mathbb{R}^{2}\mid A_{x}x\leq b_{x}\}$ and $u(k)\in\mathcal{U}=\{u\in\mathbb{R}\mid A_{u}x\leq b_{u}\}$, which are given as bounds on the angle $-0.5\leq\psi\leq 0.5$, the angular velocity $-1.5\leq\omega\leq 1.5$ and the input $-3\leq u\leq 3$. We choose the terminal penalty $\alpha_{f}=1\textup{e}6$ and as tightening values $\Delta_{i}=i\cdot 10^{-3}$ for $i=1,\dots,N$, where the prediction horizon of the PCBF problems (10) and (9) is $N=10$.
@@
In order to compute the terminal CBF $h_{f}$, we follow a similar procedure to [12, Section IV] by first computing a quadratic Lyapunov function $V(x)=x^{\top}Px$ and corresponding state feedback $u(k)=Kx(k)$ for the system (24). The Lyapunov function is computed such that the ellipse described by $V(x)\leq 1$ is the largest ellipse within the tightened state constraints $\bar{X}_{N-1}(0)$ such that $Kx\in\mathcal{U}$ according to the following semi-definite program
@@
[…]
@@
Finally, we compare the closed-loop performance of both algorithms in Figure 2. For this comparison, we train another neural network using the same data sets generated by sampling $h(x)$ by gridding the state space, where we choose a higher expressiveness of the model by using $3$ hidden layers with $16$ softplus activation functions. The model is trained in $30$ epochs using a batch size of $8$, a learning rate of $5\textup{e}-3$ and the MSE loss function, with a total computation time of approximately $25$min. The final validation MSE is given by $6\textup{e}-3$, which is two orders of magnitude lower compared to the two underfit models. As a performance input, we choose $u_{\mathrm{p}}(k)=2$, which aims to drive the system into the constraint $\psi=0.5$ and causes a slight chattering in the velocity at the constraint in closed-loop. By setting $c_{\Delta}=0.15$, we observe that in closed-loop, for 5 different initial conditions using the same initial angle $\psi=0$ and different velocities $\omega$, the system is allowed to converge much slower to $\mathcal{S}_{\mathrm{PB}}$, while still converging to the constraint $\psi=0.5$. The average solve times for the two methods over the $150$ simulated time steps are given by $0.021s$ for the PCBF algorithm in [12], which is solved using CVXPY [43] and MOSEK [44], and $0.011s$ for the approximate PCBF safety filter, which is solved using CasADi [45] and IPOPT [46]. We note that even for the small time horizon of $N=10$, we observe an almost $2$-times computational speedup. While the computation time of the PCBF algorithm is expected to increase with increasing time horizon, the online computation of the approximation based safety filter is independent of the considered time horizon and only the offline data collection time will be impacted.
@@
Figure 2: Closed-loop comparison of the PCBF-based safety filter in (10) and (9) with the proposed approximation scheme for system (24) with states given by the angle $\psi$ (top) and the angular velocity $\omega$ (middle). The explicit approximation enables tuning the rate of the PCBF decrease through the tuning parameter $c_{\Delta}\in(0,1]$, resulting in a slower convergence of the states to $\mathcal{S}_{\mathrm{PB}}$, seen by the corresponding CBF value $h(x)$ (bottom).

### v2

@@ 5.2 ISS of the Approximate PCBF Algorithm
In the proof of Theorem 5.21, the error $|e_{h}(x)|$ can also be bounded by a class $\mathcal{K}$ function $\gamma$ and a constant term $\tilde{\epsilon}_{h}\leq\epsilon_{h}$, such that $|e_{h}(x)|\leq\gamma(x)+\tilde{\epsilon}_{h}$. If it holds that $\forall x\in\mathcal{D}_{\mathrm{PB}}$, $c_{\Delta}\Delta h_{\mathrm{PB}}(x)>2\gamma(x)$, the decrease condition is still satisfied with respect to the smaller constant $2\tilde{\epsilon}_{h}+2\epsilon_{h}$. This implies that approximation errors close to $\partial\mathcal{S}_{\mathrm{PB}}$ have a bigger impact on the closed-loop stability than errors when $h_{\mathrm{PB}}$ is large, which can be leveraged when performing the approximation as discussed in Section 6.
This analysis applies to any continuous approximation of a CBF according to Definition 1 or 4. It is especially useful when formal verification methods as discussed in [51] are not applicable due to large networks and when exogenous disturbances are present. We note that for many neural network-based CBF approximations, guaranteeing that a decrease condition is satisfied is encouraged by specific loss functions during training, as discussed in [49], but cannot typically be directly enforced. If the optimal value function $h_{\mathrm{PB}}$ is Lipschitz continuous, the approach in [31] could be applied to obtain a valid approximation, however the resulting optimization problem suffers from computational complexity issues.
The established ISS result therefore provides an understanding of the closed-loop system behaviour when the CBF decrease constraint is possibly violated, either through approximation errors or disturbances. It implies that the system converges to and remains in a neighbourhood of the safe set. Additionally, sampling-based approaches as used in [25] can provide high probability guarantees for invariance of $\mathcal{D}_{\mathrm{PB}}$ given $\hat{h}_{\mathrm{PB}}$ instead of the provided error bound.
The proposed result directly implies ISS of the PCBF algorithm in [12] as it holds that $\epsilon_{h}=0$. Furthermore, if $\Delta\hat{h}_{\mathrm{PB}}(x)>0$ for all $x\in\hat{\mathcal{D}}\setminus\hat{\mathcal{S}}$, with $\hat{\mathcal{S}}=\{x\mid\hat{h}_{\mathrm{PB}}(x)\leq 0\}$ and $\hat{\mathcal{D}}=\{x\mid\hat{h}_{\mathrm{PB}}(x)\leq\hat{\gamma}\}$ for $\hat{\gamma}>0$, then $\max\{0,\hat{h}_{\mathrm{PB}}(x)\}$ directly defines a CBF according to Definition 4.
@@ 6.1 Unstable Linear System
$x(k+1)=\begin{bmatrix}1&T_{s}\\ T_{s}\frac{g}{l}&1-T_{s}c\end{bmatrix}x(k)+\begin{bmatrix}0\\ 5T_{s}\end{bmatrix}u(k),$
where the states are given by the angle and angular velocity, i.e., $x=\left[\psi\;\;\omega\right]^{\top}\in\mathbb{R}^{2}$, the input $u\in\mathbb{R}$ and the dynamics are defined through the sampling time $T_{s}=0.1$, the damping coefficient is $c=1$, the length $l=1.3$ and $g=9.81$. The system is subject to polytopic state and input constraints $x(k)\in\mathcal{X}=\{x\in\mathbb{R}^{2}\mid A_{x}x\leq b_{x}\}$ and $u(k)\in\mathcal{U}=\{u\in\mathbb{R}\mid A_{u}u\leq b_{u}\}$, which are given as bounds on the angle $0.5\leq\psi\leq 0.5$, the angular velocity $1.5\leq\omega\leq 1.5$ and the input $3\leq u\leq 3$. We choose the terminal penalty $\alpha_{f}=1$e6 and tightening values $\Delta_{i}=i\cdot 1$e-3 for $i=1,\dots,N$, with prediction horizon $N=10$.
@@
In order to compute the terminal CBF $h_{f}$, we follow a similar procedure to [12, Section IV] by first computing a quadratic Lyapunov function $V(x)=x^{\top}Px$ and corresponding state feedback $u(k)=Kx(k)$ for the system (6.1). The Lyapunov function is computed such that the ellipse described by $V(x)\leq 1$ is the largest ellipse within the tightened state constraints $\bar{X}_{N-1}(0)$ such that $Kx\in\mathcal{U}$ according to the following semi-definite program
@@
The training and validation data sets $\mathbb{D}=\{(x_{i},h(x_{i}))\}$ and $\mathbb{D}_{\textup{PB}}=\{(x_{i},h_{\mathrm{PB}}(x_{i}))\}$ are collected by gridding the state space in $2\leq\psi\leq 2$ and $8\leq\omega\leq 8$ as well as additional data points close to the constraints in $\{x\mid 0.8b_{x}\leq A_{x}x\leq 1.2b_{x}\}$ using a rejection sampler for $h_{\mathrm{PB}}(x_{i})\leq 10$, resulting in approximately $3.5\textup{e}5$ data points. The data collection for the training data of the extended PCBF took approximately 4545min without parallelization on an Intel Core i9-7940X CPU @ 3.103.10GHz. In order to illustrate the advantages of the continuous extension of the PCBF, we approximate $h_{\mathrm{PB}}$ and $h$ such that the models $\hat{h}_{\mathrm{CE}}$ and $\hat{h}_{\mathrm{PB}}$ underfit the data. In order to achieve this underfit, we train two neural networks using PyTorch [42] on $\mathbb{D}$ and $\mathbb{D}_{\textup{PB}}$ with two hidden layers of 4 neurons using the softplus activation function. The networks are trained for $50$ epochs using a learning rate of $1\textup{e}-3$, a batch size of $128$ and the mean squared error (MSE) as a loss function. Note that the expressiveness of the chosen networks is not high enough to accurately fit the data, such that after training the MSE with respect to the validation set is given by $0.25$ and $0.11$ for $\hat{h}_{\mathrm{CE}}$ and $\hat{h}_{\mathrm{PB}}$, respectively. In order to compare the performance of both approximations in the safety filter formulation, we sample $1\textup{e}4$ random states within $\mathcal{X}$. The total safety filter interventions in terms of the norm of the difference between the computed input and a performance input $u_{\mathrm{p}}(k)=0$, normalised by the interventions of the PCBF algorithm in [12], is given in Table 1. We observe that the underfit models $\hat{h}_{\mathrm{PB}}$ and $\hat{h}_{\mathrm{CE}}$ using $c_{\Delta}=1$ show a similar magnitude of safety interventions in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, being more conservative with $5\%$ and $10\%$ more interventions, respectively, compared to the full optimal control based algorithm. However, within $\mathcal{S}_{\mathrm{PB}}$, the approximation $\hat{h}_{\mathrm{PB}}$ shows a $9$ times increase in safety interventions due to learning errors when $h_{\mathrm{PB}}(x)=0$ possibly leading to positive values of $\hat{h}_{\mathrm{PB}}$, which results in modifying the proposed input. The continuous extension reduces the impact of these errors by approximating negative values in $\mathcal{S}_{\mathrm{PB}}$, such that the safety interventions are only $3.5$ times as large. As the explicit approximation enables the introduction of the additional hyperparameter $c_{\Delta}$ to tune the rate of PCBF decrease in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, we also compare the magnitude using $c_{\Delta}=0.2$. It can be seen that this results in a significant decrease in interventions in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, reducing the magnitude by a factor of $2$ for both approximations.
@@
Finally, we compare the closed-loop performance of both algorithms in Figure 2. For this comparison, we train another neural network using the same data sets generated by sampling $h(x)$ by gridding the state space, where we choose a higher expressiveness of the model by using $3$ hidden layers with $16$ softplus activation functions. The model is trained in $30$ epochs using a batch size of $8$, a learning rate of $5\textup{e}-3$ and the MSE loss function, with a total computation time of approximately 2525min. The final validation MSE is given by $6\textup{e}-3$, which is two orders of magnitude lower compared to the two underfit models. Following the arguments in [25], with a confidence of 1-11e-5, the probability that the domain $\{x\mid h_{\mathrm{PB}}(x)\leq 10\}$, where training data was collected, is robustly invariant is greater than $99.75\%$. As a performance input, we choose $u_{\mathrm{p}}(k)=2$, which aims to drive the system into the constraint $\psi=0.5$ in an adversarial sense and causes a slight chattering in the velocity at the constraint in closed-loop. In practice, such chattering can be reduced using an additional regularisation penalising the change in input as done in [1]. By setting $c_{\Delta}=0.15$, we observe that in closed-loop, for 5 different initial conditions using the same initial angle $\psi=0$ and different velocities $\omega$, the system is allowed to converge much slower to $\mathcal{S}_{\mathrm{PB}}$, while still converging to the constraint $\psi=0.5$. We additionally provide a comparison to a soft-constrained predictive safety filter (PSF) in Figure 3 for an initial velocity of $\omega=5$. The soft-constrained PSF exhibits the same behaviour as the PCBF algorithm when choosing a large slack penalty, denoted as $c_{\xi}$, in the cost and does not converge to the constraints for low values. The average solve times over the $150$ simulated time steps are given by $0.017$s for the soft-constrained PSF, $0.021$s for the PCBF algorithm in [12] and $0.011$s for the approximate PCBF safety filter, which are solved using CasADi [45] and IPOPT [46]. We note that even for the small time horizon of $N=10$, we observe a computational speedup of over $50\%$. While the computation time of the PCBF algorithm and soft-constrained PSF are expected to increase with increasing time horizon, the online computation of the approximation based safety filter is independent of the considered time horizon and only the offline data collection time will be impacted.
@@
Figure 2: Closed-loop comparison of the PCBF-based safety filter in (9) and (8) with the proposed approximation scheme for system (6.1) with states given by the angle $\psi$ (top) and the angular velocity $\omega$ (middle). The explicit approximation enables tuning the rate of the PCBF decrease through the tuning parameter $c_{\Delta}\in(0,1]$, resulting in a slower convergence of the states to $\mathcal{S}_{\mathrm{PB}}$, seen by the corresponding CBF value $h(x)$ (bottom).
Figure 3: Closed-loop comparison of soft-constrained predictive safety filters and the PCBF-based safety filter in (9) and (8) with the proposed scheme for an unstable pendulum. The angle $\psi$ (top) is depicted given a high initial angular velocity. The proposed formulation enables tuning the rate of the PCBF decrease through the tuning parameter $c_{\Delta}\in(0,1]$, resulting in a slower convergence of the states to $\mathcal{S}_{\mathrm{PB}}$ as can be observed from the PCBF values (bottom).